# A unified concept of the degree of ill-posedness for compact and non-compact linear operators in Hilbert spaces under the auspices of the spectral theorem

### v1

@@ Abstract
Covering ill-posed problems with compact and non-compact operators regarding the degree of ill-posedness is a never ending story written by many authors in the inverse problems literature. This paper tries to add a new narrative and some new facets with respect to this story under the auspices of the spectral theorem. The latter states that any self-adjoint and bounded operator is unitarily equivalent to a multiplication operator on some (semi-finite) measure space. We will exploit this fact and derive a distribution function from the corresponding multiplier, the growth behavior of which at zero allows us to characterize the degree of ill-posedness. We prove that this new concept coincides with the well-known one for compact operators (by means of their singular values), and illustrate the implications along examples including the Hausdorff moment operator and convolutions.
@@ 1 Introduction
that is ill-posed in correspondence with the fact that ${\operatorname{ran}}(T)$ is assumed to be a non-closed subset of $\mathcal{Y}$. In the seminal paper [27], the case of compact operators $T$ well-studied in the literature was called ill-posedness of type II, and their strength of ill-posedness is clearly expressed by the decay rate of the singular values of $T$. The majority of occurring linear ill-posed problems (e.g., Fredhom and Volterra integral equations of first kind over some bounded domain in $\mathbb{R}^{d}$ as well as history match problems for the heat equation with homogeneous boundary conditions) belong to that type. Such problems with compact forward operators are extensively analyzed in the inverse problems literature, see e.g. [22, 25, 28]. In the Hilbert space setting, the alternative ill-posedness of type I is characterized by non-compact operators $T$ with non-closed range and arises for convolution operators on unbounded domains (see Subsection 4.2 below) as well as for Cesàro operators (cf., e.g., [4, 5, 23]) and in the context of the Hausdorff moment problem (see Subsection 4.1 below). The absence of singular values makes the things more difficult for non-compact $T$. Therefore, it is our goal to make a suggestion for characterizing the nature of ill-posedness by means of spectral properties of the operator $T$ for both types of ill-posedness in a unified manner
@@
An illustration of those operators which are in our focus is given in Figure 1 at the end of this introduction. Let us briefly explain here this illustration of operator situations considered in the paper at hand. If $T:\mathcal{X}\to\mathcal{Y}$ is an injective, bounded linear operator, then $T^{*}T:\mathcal{X}\to\mathcal{X}$ is positive and self-adjoint and has as a consequence of Halmos’ spectral theorem from [10] a spectral decomposition with a multiplier function $\lambda$ on some measure space $\left(\Omega,\mu\right)$ (see Section 3 for details). Under the auspices of the spectral theorem, we will classify the operators $T$ by means of properties of the multiplier function $\lambda$. Our focus is on the classification of ill-posed situations that are indicated by the occurrence of essential zeros of the function $\lambda$. Such ill-posed problem are characterized in Figure 1 by two areas marked hatched in red. The right crescent of this red area is devoted to the subcase of compact operators (type II ill-posedness), whereas the left part of the red area is reserved for non-compact operators (type I ill-posedness). However, we have to exclude the subsituations $(A)$, $(B)$ and $(C)$, marked in black. First the exclave area (A) in black expresses the case of non-compact operators, where the measure $\mu(\Omega)$ is finite (Assumption 1(a) below is violated). This case is widely discussed in the literature, in particular for multiplication operators on $\mathcal{X}=L^{2}(0,1)$ with Lebesgue measure $\mu=\mu_{\mathrm{Leb}}$, and we give an overview of literature and specific approaches for that subcase in the appendix. The second black half-circle (B) in Figure 1 collects exceptional situations, where the corresponding distribution function $\Phi$ (see formula (7)) is not informative and Assumption 1(b) below is violated. Such situation is made explicit in Example 5 below. It also occurred in [16, Example 1] in form of a diagonal operator in $\ell^{2}$, where only a subsequence of the diagonal elements tends to zero. The last small black semi-circle (C) is reserved for possibly occurring situations, where $\mu$ cannot chosen to be the Lebesgue measure $\mu_{\mathrm{Leb}}$ on some subset of $\mathbb{R}^{d}$ (Assumption 2 is violated). All three cases marked in black are beyond our focus of this paper.
@@
The paper is organized as follows: In the following Section 2 we recall the well-known concept of the degree of ill-posedness for compact operators and derive equivalent characterizations of the important interval of ill-posedness. Section 3 is devoted to the non-compact case. We there state the Halmos version of the spectral theorem to be used in the following, define the corresponding distribution function, and illustrate along examples how it can be used to define the degree of ill-posedness. Afterwards, we give our definition and relate it to the compact subcase. Since the focus of our studies is on multiplication operators on infinite measure spaces, a glimpse of the finite measure case is given in an appendix. In Section 4 we discuss examples such as the Hausdorff moment operator and convolutions and derive their degree of ill-posedness. Section 5 is briefly devoted to the case of unbounded operators, which in principle could be handled similarly. We end our manuscript by a brief conclusion and outlook in Section 6.
2 The degree of ill-posedness for compact operators
@@
The ill-posedness of $T$ is due to the fact that $\lim_{n\to\infty}\sigma_{n}=0$, and the decay rate of $\{\sigma_{n}\}_{n\in\mathbb{N}}$ indicates the severeness of this ill-posedness. The first case of a polynomial decay of singular values for compact operators is well-known in the literature as moderate ill-posedness (cf., e.g., [12])), the second case of an exponential decay as severe ill-posedness (cf., e.g., [6]). The third case of a logarithmic decay can consequently be characterized as mild ill-posedness (cf., e.g, [1]).
@@
Let $T:\mathcal{X}\to\mathcal{Y}$ be an injective and compact linear operator between infinite-dimensional Hilbert spaces $\mathcal{X}$ and $\mathcal{Y}$, and let $\sigma_{n}$ be the corresponding singular values. We define the interval of ill-posedness of $T$ as
@@
The interval of ill-posedness of $T$ is given by
@@
Let $T:\mathcal{X}\to\mathcal{Y}$ be an injective and compact linear operator between infinite-dimensional Hilbert spaces $\mathcal{X}$ and $\mathcal{Y}$, and let $\Phi$ its counting function as in (3). Then the interval of ill-posedness of $T$ is given as
@@
The above theorem and the derived corollary show that we can express the degree of ill-posedness in the compact case equivalently by means of the counting function $\Phi$ from (3). In the following, we will use this as a blue-print to consistently extend the corresponding definition to the non-compact case under certain restrictions.
@@
For Example 3, solving the operator equation (1) requires to find $u=(-\Delta)g$, which means that second partial derivatives of $g$ have to be calculated. The factor $2$ in the degree of ill-posedness $s=2/d$ seems to express this. Namely, for the compact embedding operator $T:H^{p}((0,1)^{d})\to L^{2}((0,1)^{d})$ from the Sobolev space of degree $p>0$ to $L^{2}$ in $d$ dimensions we have singular values as $\sigma_{n}\asymp n^{-\frac{p}{d}}$ (cf.,e.g.,[24, §3c]) implying a degree of ill-posedness as $s=p/d$, and just for $p=2$ calculations up to second partial derivatives occur.
@@ 3.1 The spectral theorem according to Halmos
Moreover, we recall the spectral theorem in multiplication operator form by Halmos [10] and take into account the additional notes in [9, p. 47] and [35, Chap. VII.1]. Then the following holds true for the normal operator $T^{*}T:\mathcal{X}\to\mathcal{X}$: There exists a locally compact topological space $\Omega$, a semi-finite (Radon) measure $\mu$ on $\Omega$ with finite or infinite measure value $\mu(\Omega)$, a real-valued function $\lambda\in L^{\infty}\left(\Omega,\mu\right)$, and a unitary mapping $W:L^{2}\left(\Omega,\mu\right)\to\mathcal{X}$ such that
@@
coincides with the spectrum ${\rm spec}(T^{*}T)$ of the operator $T^{*}T$ (cf., e.g., [9, Theorem 2.1(g)]) and is hence a subset of $[0,\|T^{*}T\|]=[0,\|T\|^{2}]$. Let us mention that the injectivity requirement for $T$ made throughout this paper implies that $\mu(\{\omega\in\Omega:\,\lambda(\omega)=0\})=0$.
As a consequence of the above mentioned coincidence between spectrum of $T^{*}T$ and the essential range of $\lambda$, the operator equation (1) is ill-posed if and only if zero belongs to the essential range of the multiplier function $\lambda$. Our goal is to characterize the ill-posedness of the original operator equation (1) and its associated normal equation by properties of this multiplier function $\lambda$ in (6) and their implications. However, this is not immediately possible without further restrictions, since the spectral decomposition (5) is not uniquely determined for the operator $T$ as the following lemma indicates. Note that this observation was already employed in [3]:
@@
The above non-uniqueness will lead to severe problems when defining the degree of ill-posedness for non-compact $T$. To illustrate this, recall that our definition of the degree of ill-posedness for compact operators via the counting function $\Phi$ from (3) explicitly relies on the counting measure $\mu_{disc}=\#$ in (3). If we allow for other measures, we will obtain completely different asymptotical behaviors of $\Phi$. As an example we next consider a history match problem.
@@
with a given terminal time $\bar{t}>0$, and denote by $T_{\mathrm{BHE}}:L^{2}\left(\pi,\pi\right)\to L^{2}\left(-\pi,\pi\right)$. The corresponding operator equation (1) is widely understood to be exponentially ill-posed, and this is reflected by the series representation
@@
This indicates a completely different degree of ill-posedness, and hence the counting measure has to be fixed to obtain a meaningful concept.
@@
3.2 The distribution function
Inspired by the counting function (3) we will consider, as an analog under the auspices of the spectral theorem, the non-negative, non-increasing and right-continuous distribution function
@@ 3.2 The distribution function
3.3 Ill-posedness of multiplication operators on the real axis
To enter the world of ill-posed linear operator equations (1) with injective, bounded and non-compact operators $T:\mathcal{X}\to\mathcal{Y}$, we first consider pure multiplication operators for $\mathcal{X}=\mathcal{Y}=L^{2}(\Omega)$ on the real axis with a measure space $(\Omega,\mu_{\mathrm{Leb}})$, where $\Omega=\mathbb{R}=(-\infty,+\infty)$ is combined with the Lebesgue measure $\mu_{\mathrm{Leb}}$ over $\mathbb{R}$.
In this section, our study restricts to the case of multiplication operators $M_{\lambda}=T^{*}T$ defined as
$[M_{\lambda}\,x](\omega):=\lambda(\omega)\cdot x(\omega)\qquad(\mu-\mbox{a.e.~{}on}\;\Omega)$ (8)
with multiplier functions $\lambda$ in (8), which are continuous everywhere on $\mathbb{R}$ and obey the limit condition $\lim_{\omega\to\pm\infty}\lambda(\omega)=0$. Then we always have a non-compact operator $T^{*}T$ with purely continuous spectrum ${\rm spec}(T^{*}T)=[0,\|T^{*}T\|]={\rm essran}(\lambda)=[0,\max\limits_{\omega\in\mathbb{R}}\lambda(\omega)]$ that implies ill-posedness of type I in the sense of Nashed of the associated operator equation (1). As a first attempt to generalize the degree of ill-posedness, and inspired by Corollary 1, we will consider the interval
$\left[A_{\lambda,\mu_{\mathrm{Leb}}},B_{\lambda,\mu_{\mathrm{Leb}}}\right]:=\left[\liminf\limits_{\epsilon\searrow 0}\frac{\log(\epsilon)}{-2\log\left(\Phi_{\lambda,\mu_{\mathrm{Leb}}}(\epsilon)\right)},\limsup\limits_{\epsilon\searrow 0}\frac{\log(\epsilon)}{-2\log\left(\Phi_{\lambda,\mu_{\mathrm{Leb}}}(\epsilon)\right)}\right]$ (9)
derived from the distribution function $\Phi_{\lambda,\mu_{\mathrm{Leb}}}$ in (7). As the variants in Example 7 below will show, also here moderately, severely and mildly ill-posed problems can be distinguished intuitively.

### v2

# A unified concept of the degree of ill-posedness for compact and non-compact linear operator equations in Hilbert spaces under the auspices of the spectral theorem

@@ Abstract
Covering ill-posed problems with compact and non-compact operators regarding the degree of ill-posedness is a never ending story written by many authors in the inverse problems literature. This paper tries to add a new narrative and some new facets with respect to this story under the auspices of the spectral theorem. The latter states that any self-adjoint and bounded operator is unitarily equivalent to a multiplication operator on some (semi-finite) measure space. We will exploit this fact and derive a distribution function from the corresponding multiplier, the growth behavior of which at zero allows us to characterize the degree of ill-posedness. We prove that this new concept coincides with the well-known one for equations with compact operators (by means of their singular values), and illustrate the implications along examples with non-compact operators, including the Hausdorff moment operator and convolutions.
@@ 1 Introduction
that is ill-posed in correspondence with the fact that ${\operatorname{ran}}(T)$ is assumed to be a non-closed subset of $\mathcal{Y}$. In the seminal paper [29], the case of compact operators $T$ well-studied in the literature was called ill-posedness of type II, and their strength of ill-posedness is clearly expressed by the decay rate of the singular values of $T$. The majority of occurring linear ill-posed problems (e.g., Fredhom and Volterra integral equations of first kind over some bounded domain in $\mathbb{R}^{d}$ as well as history match problems for the heat equation with homogeneous boundary conditions) belong to that type. Such problems with compact forward operators are extensively analyzed in the inverse problems literature, see e.g. [23, 26, 30]. In the Hilbert space setting, the alternative ill-posedness of type I is characterized by non-compact operators $T$ with non-closed range and arises for convolution operators on unbounded domains (see Subsection 4.2 below) as well as for Cesàro operators (cf., e.g., [4, 5, 24]) and in the context of the Hausdorff moment problem (see Subsection 4.1 below). The absence of singular values makes the things more difficult for non-compact $T$. Therefore, it is our goal to make a suggestion for classifying the strength of ill-posedness also for non-compact operators $T$ with non-closed range by means of their spectral properties.
Covering ill-posed problems with compact and non-compact operators regarding the nature and degree of ill-posedness is a never ending story written by many authors in the inverse problems literature. For example, we refer to the claim of Nashed in [29] that equations with non-degenerating compact operators mapping between infinite dimensional Hilbert spaces are always more ill-posed than those with non-compact ones. Arguments for that claim come from the comparison of ill-posed situations with two linear operators by partial orderings and associated range inclusions (see, e.g, [17, 27]). In this context, the range of a non-compact operator can never be a subset of the range of a compact operator. Also in [17], however, it was shown by observing the modulus of injectivity for various discretizations that the ill-posedness nature of non-compact operators is completely different from the nature of compact ones and that hence the simple claim of Nashed is rather problematic, because sequences of compact operators do never converge in norm to a non-compact operator. In the present paper, we try to discuss some new facets of the story of covering compact and non-compact cases under the auspices of the spectral theorem.
@@
An illustration of those operators which are in our focus is given in Figure 1 at the end of this introduction. Let us briefly explain here this illustration of operator situations considered in the paper at hand. If $T:\mathcal{X}\to\mathcal{Y}$ is an injective, bounded linear operator, then $T^{*}T:\mathcal{X}\to\mathcal{X}$ is positive and self-adjoint and has as a consequence of Halmos’ spectral theorem from [11] a spectral decomposition with a multiplier function $\lambda$ on some measure space $\left(\Omega,\mu\right)$ (see Section 3 for details). Under the auspices of the spectral theorem, we will classify the operators $T$ by means of properties of the multiplier function $\lambda$. Our focus is on the classification of ill-posed situations that are indicated by the occurrence of essential zeros of the function $\lambda$. Such ill-posed problem are characterized in Figure 1 by two areas marked hatched in red. The right crescent of this red area is devoted to the subcase of compact operators (type II ill-posedness), whereas the left part of the red area is reserved for non-compact operators (type I ill-posedness). However, we have to exclude the subsituations $(A)$, $(B)$ and $(C)$, marked in black. First the exclave area (A) in black expresses the case of non-compact operators, where the measure $\mu(\Omega)$ is finite (Assumption 1(a) below is violated). This case is widely discussed in the literature, in particular for multiplication operators on $\mathcal{X}=L^{2}(0,1)$ with Lebesgue measure $\mu=\mu_{\mathrm{Leb}}$, and we give an overview of literature and specific approaches for that subcase in the appendix. The second black half-circle (B) in Figure 1 collects exceptional situations, where the corresponding distribution function $\Phi$ (see formula (7)) is not informative and Assumption 1(b) below is violated. Such situation is made explicit in Example 5 below. It also occurred in [17, Example 1] in form of a diagonal operator in $\ell^{2}$, where only a subsequence of the diagonal elements tends to zero. The last small black semi-circle (C) is reserved for possibly occurring situations, where $\mu$ cannot chosen to be the Lebesgue measure $\mu_{\mathrm{Leb}}$ on some subset of $\mathbb{R}^{d}$ (Assumption 2 is violated). All three cases marked in black are beyond our focus of this paper.
@@
The paper is organized as follows: In the following Section 2 we recall the well-known concept of the degree of ill-posedness for equations with compact operators and derive equivalent characterizations of the important interval of ill-posedness. Section 3 is dedicated to the inclusion of the non-compact case. We there state the Halmos version of the spectral theorem to be used in the following, define the corresponding distribution function, and illustrate along examples how it can be used to define the degree of ill-posedness. Afterwards, we give our definition and relate it to the compact subcase. Since the focus of our studies is on multiplication operators on infinite measure spaces, a glimpse of the finite measure case is given in an appendix. In Section 4 we discuss prominent examples of non-compact operators with non-closed range, such as the Hausdorff moment operator and convolutions, and derive their degree of ill-posedness. Section 5 is briefly devoted to the case of unbounded operators, which in principle could be handled similarly. We end our manuscript by a brief conclusion and outlook in Section 6.
2 The degree of ill-posedness for equations with compact operators
@@
The ill-posedness of equations involving $T$ is due to the fact that $\lim_{n\to\infty}\sigma_{n}=0$, and the decay rate of $\{\sigma_{n}\}_{n\in\mathbb{N}}$ indicates the severeness of this ill-posedness. The first case of a polynomial decay of singular values for compact operators is well-known in the literature as moderate ill-posedness (cf., e.g., [13])), the second case of an exponential decay as severe ill-posedness (cf., e.g., [6]). The third case of a logarithmic decay can consequently be characterized as mild ill-posedness (cf., e.g, [1]).
@@
Let $T:\mathcal{X}\to\mathcal{Y}$ be an injective and compact linear operator between infinite-dimensional Hilbert spaces $\mathcal{X}$ and $\mathcal{Y}$, and let $\sigma_{n}$ be the corresponding singular values. We define the interval of ill-posedness of the operator equation (1) as
@@
The interval of ill-posedness of (1) is given by
@@
Let $T:\mathcal{X}\to\mathcal{Y}$ be an injective and compact linear operator between infinite-dimensional Hilbert spaces $\mathcal{X}$ and $\mathcal{Y}$, and let $\Phi$ its counting function as in (3). Then the interval of ill-posedness of (1) is given as
@@
The above theorem and the derived corollary show that we can express the degree of ill-posedness in the compact case equivalently by means of the counting function $\Phi$ from (3). In the following, we will use this as a blueprint to consistently extend the corresponding definition to the non-compact case under certain restrictions.
@@
For Example 3, solving the operator equation (1) requires to find $u=(-\Delta)g$, which means that second partial derivatives of $g$ have to be calculated. The factor $2$ in the degree of ill-posedness $s=2/d$ seems to express this. Namely, for the compact embedding operator $T:H^{p}((0,1)^{d})\to L^{2}((0,1)^{d})$ from the Sobolev space of degree $p>0$ to $L^{2}$ in $d$ dimensions we have singular values as $\sigma_{n}\asymp n^{-\frac{p}{d}}$ (cf.,e.g.,[25, §3c]) implying a degree of ill-posedness as $s=p/d$, and just for $p=2$ calculations up to second partial derivatives occur.
@@ 3.1 The spectral theorem according to Halmos
Moreover, we recall the spectral theorem in multiplication operator form by Halmos [11] and take into account the additional notes in [10, p. 47] and [37, Chap. VII.1]. Then the following holds true for the normal operator $T^{*}T:\mathcal{X}\to\mathcal{X}$: There exists a locally compact topological space $\Omega$, a semi-finite (Radon) measure $\mu$ on $\Omega$ with finite or infinite measure value $\mu(\Omega)$, a real-valued function $\lambda\in L^{\infty}\left(\Omega,\mu\right)$, and a unitary mapping $W:L^{2}\left(\Omega,\mu\right)\to\mathcal{X}$ such that
@@
coincides with the spectrum ${\rm spec}(T^{*}T)$ of the operator $T^{*}T$ (cf., e.g., [10, Theorem 2.1(g)]) and is hence a subset of $[0,\|T^{*}T\|]=[0,\|T\|^{2}]$. Let us mention that the injectivity requirement for $T$ made throughout this paper implies that $\mu(\{\omega\in\Omega:\,\lambda(\omega)=0\})=0$.
The operator equation (1) is ill-posed if and only if $0\in{\rm spec}(T^{*}T)$. Then as a consequence of the above mentioned coincidence between spectrum of $T^{*}T$ and the essential range of the multiplier function $\lambda$, the criterion of ill-posedness for (1) is the occurrence of an essential zero of the function $\lambda$, which means that zero belongs to the essential range of $\lambda$. Our goal is to characterize the ill-posedness of the original operator equation (1) and its associated normal equation by properties of this multiplier function $\lambda$ in (6) and their implications. However, this is not immediately possible without further restrictions, since the spectral decomposition (5) is not uniquely determined for the operator $T$ as the following lemma indicates. Note that this observation was already employed in [3]:
@@
The above non-uniqueness will lead to severe problems when defining the degree of ill-posedness for equations with non-compact $T$, but also in the compact case it brings in a subjective factor. To illustrate this, recall that our definition of the degree of ill-posedness for compact operators via the counting function $\Phi$ from (3) explicitly relies on the counting measure $\mu_{disc}=\#$ in (3). If we allow for other measures, we will obtain completely different asymptotical behaviors of $\Phi$. As an example we next consider a history match problem.
@@
with a given terminal time $\bar{t}>0$, and denote by $T_{\mathrm{BHE}}:L^{2}\left({-\pi},\pi\right)\to L^{2}\left(-\pi,\pi\right)$. The corresponding operator equation (1) is widely understood to be exponentially ill-posed, and this is reflected by the series representation
@@
This indicates a completely different degree of ill-posedness, and hence the counting measure has to be fixed as benchmark measure for compact operators to obtain a meaningful concept.
@@
3.2 The distribution function measuring the strength of ill-posedness
Inspired by the counting function (3) we will consider, as an analog under the auspices of the spectral theorem and now applicable to both compact and non-compact operators $T$, the non-negative, non-increasing and right-continuous distribution function
@@ 3.2 The distribution function measuring the strength of ill-posedness
In case that Assumption 1 holds true, it makes sense to formulate an analog to Definition 1 under the auspices of the spectral theorem as follows:
Definition 2.
Let $\mathcal{X}$ and $\mathcal{Y}$ denote infinite-dimensional Hilbert spaces. Moreover, let $T:\mathcal{X}\to\mathcal{Y}$ be an injective and bounded linear operator with non-closed range, each mapping between $\mathcal{X}$ and $\mathcal{Y}$ such that the Assumption 1 is satisfied w.r.t. the distribution function $\Phi_{\lambda,\mu}$ from (7). Then the interval of ill-posedness of the operator equation (1) is defined as
$\left[A_{\Phi_{\lambda,\mu}},B_{\Phi_{\lambda,\mu}}\right]:=\left[\liminf\limits_{\epsilon\searrow 0}\frac{\log(\epsilon)}{-2\log\left(\Phi_{\lambda,\mu}(\epsilon)\right)},\limsup\limits_{\epsilon\searrow 0}\frac{\log(\epsilon)}{-2\log\left(\Phi_{\lambda,\mu}(\epsilon)\right)}\right],$ (8)
and the operator equation (1) is called mildly ill-posed whenever $A_{\Phi_{\lambda,\mu}}=B_{\Phi_{\lambda,\mu}}=0$, severely ill-posed whenever $A_{\Phi_{\lambda,\mu}}=B_{\Phi_{\lambda,\mu}}=\infty$, moderately ill-posed whenever $0<A_{\Phi_{\lambda,\mu}}\leq B_{\Phi_{\lambda,\mu}}<\infty$, and ill-posed of degree $s>0$ if $A_{\Phi_{\lambda,\mu}}=B_{\Phi_{\lambda,\mu}}=s$. Then the distribution functions
Theorem 2.
For compact operators $T$, the intervals and degrees of ill-posedness from Definitions 1 and 2 agree.
Proof.
For compact operators $T$ with singular values $\{\sigma_{n}\}_{n=1}^{\infty}$ there is a unitary operator $W:L^{2}(\mathbb{N},\mu_{disc})\to\mathcal{X}$ for $\Omega=\mathbb{N}$ and the counting measure $\mu=\mu_{disc}$ such that the formulas (5) and (6) hold with the multiplier function $\lambda(n)=\sigma_{n}^{2}\;(n\in\mathbb{N})$. Then the counting function $\Phi$ from (3) and the distribution function $\Phi_{\lambda,\mu}$ from (7) match. Taking into account the result of Corollary 1, this proves the theorem. ∎
[…]
3.3 Ill-posedness of equations with multiplication operators on the real axis
To enter the world of ill-posed linear operator equations (1) for injective, bounded and non-compact operators $T:\mathcal{X}\to\mathcal{Y}$ with non-closed range, we first consider in this subsection pure multiplication operators $T:=M_{\lambda}$ defined as $[M_{\lambda}\,\xi](\omega):=\lambda(\omega)\cdot\xi(\omega)\;(\mu-\mbox{a.e.~{}on}\;\Omega)$ for $\mathcal{X}=\mathcal{Y}=L^{2}(\Omega)$ on the real axis with a measure space $(\Omega,\mu_{\mathrm{Leb}})$, where $\Omega=\mathbb{R}=(-\infty,+\infty)$ is combined with the Lebesgue measure $\mu_{\mathrm{Leb}}$ over $\mathbb{R}$. We assume that the multiplier functions $\lambda$ is continuous everywhere on $\mathbb{R}$ and obeys the limit condition $\lim_{\omega\to\pm\infty}\lambda(\omega)=0$. Then we always have a non-compact operator $T^{*}T$ with purely continuous spectrum ${\rm spec}(T^{*}T)=[0,\|T^{*}T\|]={\rm essran}(\lambda)=[0,\max\limits_{\omega\in\mathbb{R}}\lambda(\omega)]$ that implies ill-posedness of type I in the sense of Nashed of the associated operator equation (1). Definition 2 now applies here for this class of non-compact operators with non-closed range by considering the interval of ill-posedness (8) derived from the distribution function (7) with $\mu=\mu_{Leb}$. As the variants in Example 7 below will show, also here moderately, severely and mildly ill-posed problems can be distinguished intuitively.